OpenStudy (anonymous):

What is the solution to the proportion? n/18 = 5/9 A. n = 2.5 B. n = 5 C. n = 10 D. n = 15

OpenStudy (anonymous):

@Shay17

OpenStudy (campbell_st):

multiply both sides of the equation by 18 and you get \[n = \frac{18 \times 5}{9}\]

OpenStudy (anonymous):

if i guess its b

OpenStudy (campbell_st):

lol.. wow 5/18 = 5/9

OpenStudy (anonymous):

ok ok dont be mean